C–H HYDROGEN BONDING INTERACTIONS IN THE CARBONYL SULFIDE–FLUOROFORM DIMER

\underline{MICHAL M. SERAFIN}, SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Ave., Charleston, IL 61920 USA.

The rotational spectrum of the weakly bound dimer of carbonyl sulfide (OCS) and fluoroform (HCF_{3}) has been assigned using Fourier-transform microwave spectroscopy. The rotational constants for the normal isotopomer are A = 4745.7157(25) MHz, B = 813.9222(27) MHz, and C = 790.9238(27) MHz, and dipole moment components are μ_{a} = 0.828(4) D and μ_{b} = 0.858(7) D. The experimental rotational constants are found to be in very good agreement with ab initio optimizations at the MP2/6-311++G(2d,2p) level of calculation. This C_{s} symmetry structure has the OCS and HCF_{3} aligned such that there is a C–H...O and C–F...C interaction, with an approximate C...C intermolecular distance of 3.60 Å.